Solve by substitution or elimination method:
3x - 2y = 8
-12x + 8y = 32
Solve by substitution or elimination method:
7x - 5y = 14
-4x + y = 27
Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15
: Solve by substitution or elimination method:
3x - 2y = 8
-12x + 8y = 32
Solve by substitution or elimination method:
7x - 5y = 14
-4x + y = 27
Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15

Answer by jim_thompson5910(9921)

# 1



Start with the given system of equations:
system(3x-2y=8,-12x+8y=32)


4(3x-2y)=4(8) Multiply the both sides of the first equation by 4.


12x-8y=32 Distribute and multiply.


So we have the new system of equations:
system(12x-8y=32,-12x+8y=32)


Now add the equations together. You can do this by simply adding the two left sides and the two right sides separately like this:


(12x-8y)+(-12x+8y)=(32)+(32)


(12x+-12x)+(-8y+8y)=32+32 Group like terms.


0x+0y=64 Combine like terms. Notice how the x terms cancel out.


0=64Simplify.


Since 0=64 is NEVER true, this means that there are no solutions. So the system is inconsistent.







# 2


Note: I've made the first equation -4x+y=27 and the second equation 7x-5y=14


Start with the given system of equations:

system(-4x+y=27,7x-5y=14)



Now in order to solve this system by using substitution, we need to solve (or isolate) one variable. I'm going to solve for y.




So let's isolate y in the first equation

-4x+y=27 Start with the first equation


y=27+4x Add 4x to both sides


y=+4x+27 Rearrange the equation


y=(+4x+27)/(1) Divide both sides by 1


y=((+4)/(1))x+(27)/(1) Break up the fraction


y=4x+27 Reduce



---------------------

Since y=4x+27, we can now replace each y in the second equation with 4x+27 to solve for x



7x-5highlight((4x+27))=14 Plug in y=4x+27 into the second equation. In other words, replace each y with 4x+27. Notice we've eliminated the y variables. So we now have a simple equation with one unknown.



7x+(-5)(4)x+(-5)(27)=14 Distribute -5 to 4x+27


7x-20x-135=14 Multiply


-13x-135=14 Combine like terms on the left side


-13x=14+135Add 135 to both sides


-13x=149 Combine like terms on the right side


x=(149)/(-13) Divide both sides by -13 to isolate x



x=-149/13 Reduce





-----------------First Answer------------------------------


So the first part of our answer is: x=-149/13









Since we know that x=-149/13 we can plug it into the equation y=4x+27 (remember we previously solved for y in the first equation).



y=4x+27 Start with the equation where y was previously isolated.


y=4(-149/13)+27 Plug in x=-149/13


y=-596/13+27 Multiply


y=-245/13 Combine like terms (note: if you need help with fractions, check out this solver)



-----------------Second Answer------------------------------


So the second part of our answer is: y=-245/13









-----------------Summary------------------------------

So our answers are:

x=-149/13 and y=-245/13

which form the ordered pair






# 3




Start with the given system of equations:
system(-4x+3y=5,12x-9y=-15)


3(-4x+3y)=3(5) Multiply the both sides of the first equation by 3.


-12x+9y=15 Distribute and multiply.


So we have the new system of equations:
system(-12x+9y=15,12x-9y=-15)


Now add the equations together. You can do this by simply adding the two left sides and the two right sides separately like this:


(-12x+9y)+(12x-9y)=(15)+(-15)


(-12x+12x)+(9y+-9y)=15+-15 Group like terms.


0x+0y=0 Combine like terms. Notice how the x terms cancel out.


0=0Simplify.


Since 0=0 is ALWAYS true, this means that there are an infinite number of solutions. So the system is consistent and dependent.


Question 170002: Solve by substitution or elimination method:
3x - 2y = 8
-12x + 8y = 32
Solve by substitution or elimination method:
7x - 5y = 14
-4x + y = 27
Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15
: Solve by substitution or elimination method:
3x - 2y = 8
-12x + 8y = 32
Solve by substitution or elimination method:
7x - 5y = 14
-4x + y = 27
Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15

Answer by Electrified_Levi(95)

Hi, Hope I can help,
.
Solve by substitution or elimination method:
.
3x - 2y = 8
-12x + 8y = 32
.
Solve by substitution or elimination method:
.
7x - 5y = 14
-4x + y = 27
.
Solve by substitution or elimination method:
.
-4x + 3y = 5
12x - 9y = -15
.
First we will solve the first system with substitution
.
3x - 2y = 8
-12x + 8y = 32
.
First we need to solve for a variable in one of the two equations, doesn't matter which letter, or equation, we will solve for "y" in the first equation
.
 3x - 2y = 8 , we will move (-2y) to the right side
.
 3x - 2y = 8 =  3x - 2y + 2y = 8 + 2y =  3x = 8 + 2y , now we need to move "8" to the left side
.
 3x = 8 + 2y =  3x - 8 = 8 - 8 + 2y =  3x - 8 = 2y , to find "y" we need to divide each side by "2"
.
 3x - 8 = 2y =  (3x - 8)/2 = 2y/2 =  (3x - 8)/2 = y
.
 y = (3x - 8)/2 , since "y" is equal to  (3x - 8)/2 , we can replace "y" in the other equation with  (3x - 8)/2 , then just solve for "x"
.
 -12x + 8y = 32 =  -12x + 8((3x - 8)/2) = 32 , now just solve for "x"
.
 -12x + 8((3x - 8)/2) = 32 =  -12x + 4((3x - 8)/1) = 32 =  -12x + 4(3x - 8) = 32 , now we will use the distribution method
.
 -12x + highlight(4)(highlight(3x) - 8) = 32 =  -12x + highlight(4)(3x - highlight(8)) = 32
.
Remember the + and - signs,  -12x + 12x - 32 = 32 , adding the "x"'s
.
 -12x + 12x - 32 = 32 =  0x - 32 = 32 =  - 32 = 32 ( False )
.
This means ( when the x's or y's cancel out, and there is a false statement) that there are no solutions, these lines are parallel, there is no intersection, and therefore no solutions
.
Here is the graph of this system
.
 graph ( 300,300,-10,10,-10,10, (3x - 8)/2, (3x+8)/2 )
.
Now we will solve the second system, by elimination
.
7x - 5y = 14
-4x + y = 27
.
Elimination is when you add the two equations together, and it gets rid of a variable, first we need to make sure the x's or y's in each equation are the same, or the negative of the other
.
We can eliminate any variable ( either x or y )
.
We will get rid of the y's
.
We need to either get both y's to "5y" or "-5y" or we need to get them to "y" or "-y", we will change the second equation to "5y"
.
 -4x + y = 27 , to get "y" to change to "5y" we need to multiply each side by (5)
.
 -4x + y = 27 =  5(-4x + y) = 5(27) =  5(-4x + y) = 135
.
We will need to use the distribution method
.
 5(-4x + y) = 135 =  highlight(5)(highlight(-4x) + y) = 135 =  highlight(5)(-4x + highlight(y)) = 135
.
Remember the signs,  -20x + 5y = 135 , this is our new equation
.
Now we will bring the firt equation to our second new equation
.
 7x - 5y = 14
.
 -20x + 5y = 135
.
We will now add the equations
.
 7x - 5y = 14
.
 -20x + 5y = 135
.
 7x + (-20x) =  -13x
.
 -5y + 5y =  0y =  0
.
 14 + 135 = 149
.
It will become  -13x + 0 = 149 =  -13x = 149 , to find "x" we need to divide each side by  -13
.
 -13x = 149 =  -13x/-13 = 149/-13 =  x = 149/-13 =  x = -149/13 , we can now replace "x" with   -149/13 , in one of the two original equations
.
7x - 5y = 14
-4x + y = 27
.
We will use the second equation
.
 -4x + y = 27 =  -4(-149/13) + y = 27 =  (-4/1)(-149/13) + y = 27 =  (596/13) + y = 27
.
Now we need to move  596/13 to the right side ( we will convert "27" into "13ths"
.
 596/13 + y = 27 =  596/13 + y = 351/13 =  (596/13) - (596/13) + y = (351/13)-(596/13) =  y = (351/13)-(596/13) =  y = -245/13
.
 y = -245/13 , we can check our answers by replacing "x" and "y" in both original equations
.
 x = -149/13
.
 y = -245/13
.
First equation,  7x - 5y = 14 =  7(-149/13) - 5(-245/13) = 14 =  (7/1)(-149/13) - ((5/1)(-245/13)) = 14 =  (-1043/13) - (-1225/13) = 14 =  (-1043/13) + 1225/13 = 14 =  182/13 = 14 =  14 = 14 ( True )
.
Second equation,  -4x + y = 27 =  -4(-149/13) + (-245/13) = 27 =  (-4/1)(-149/13) - (245/13) = 27 =  (596/13) - (245/13) = 27 =  351/13 = 27 =  27 = 27 , ( True )
.
 x = -149/13
.
 y = -245/13
.
Solution sets are in the form (x,y), our solution set is (  -149/13 ,  -245/13 )
.
The graph of the system is
.
 drawing ( 500,500,-20,20,-20,20,red (circle (-149/13,-245/13,0.1)),blue (circle (-149/13,-245/13,0.2)),graph ( 500,500,-20,20,-20,20, (7x-14)/5, 4x+27))
.
The intersection is your answer
.
Now we can do the last system
.
-4x + 3y = 5
12x - 9y = -15
.
We will use the elimination method of solving this problem
.
Remember elimination is where you get rid of a variable
.
We will multiply the first equation by "3" to get rid of the "y"
.
 -4x + 3y = 5 =  3(-4x + 3y) = 3(5) =  3(-4x + 3y) = 15 we will use distribution
.
 3(-4x + 3y) = 15 =  highlight (3)(highlight(-4x) + 3y) = 15 =  highlight (3)(-4x + highlight(3y)) = 15
.
Remember signs,  -12x + 9y = 15 , now we can add the new first equation to the second equation
.
 -12x + 9y = 15
.
 12x - 9y = -15
.
 -12x + 12x =  0x =  0
.
 9y + (-9y) =  9y - 9y =  0y =  0
.
 15 + (-15) =  15 - 15 =  0
.
It will become  0 + 0 = 0 =  0 = 0 (True)
.
This means ( when both x and y cancel out, and there is a true statement) that this is the same line, both equation will be one line, there is an infinite number of solutions ( since the equations are equal to one line )( or the second line is put on the first line )
.
Here is the graph of this system
.
 graph ( 300,300, -10,10,-10,10, (4x + 5)/3, (4x+5)/3 )
.
Your answers are
.
First system of equations = " no solutions " ( since they are parallel lines )
.
Second system of equations = (  -149/13 ,  -245/13 ) ( the lines intersect at point (  -149/13 ,  -245/13 )
.
Third system of equations = " infinite solutions " ( since the two equations are the same line )
.
Hope I helped, Levi